40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log, Page 122 of 339  Not logged in ELOG logo
ID Date Author Type Category Subject
  10934   Fri Jan 23 10:07:15 2015 SteveUpdateSUSSUS Drift ETMX vs ETMY

ETMX YAW stopped drifting Jan 8, 2015

Quote:

I made little scripts to go with the sus driftmon buttons, that will servo the alignment sliders until the susyaw and suspit values match the references on the driftmon screen.

Attachment 1: SUSdrift30d.png
SUSdrift30d.png
  10933   Fri Jan 23 02:11:40 2015 JenneUpdateLSCCARM filters modified slightly

[Jenne, Diego]

One of tonight's goals was to tweak the CARM filters, so that we could engage the lowpass filter, to avoid the detuned double cavity pole resonance disturbing the CARM loop.

I increased the Q of the zeros in the FM3 boost so that it eats fewer than the original 18 degrees of phase at 100 Hz.  We kept losing lock though, so for each lock I backed off on the Q a little bit.  In the end, the filter eats 9 degrees of phase at 100 Hz.  I also moved the lowpass from 700 Hz to 1kHz, although that doesn't change the phase at 100 Hz very much.

We modified the carm_up script re: PRMI locking a little bit.  The PRMI is not so enthusiastic about locking immediately at 25% MICH fringe, so I backed that off.  It now acquires lock at a few percent, and then ramps up the offset.  Also, the MICH FM6 bounce roll filter is now turned on after lock is acquired, effectively giving it an extra second or two of delay beyond the rest of the filters.

We were able several times to get to some MICH offset and turn on the lowpass filter, but starting to reduce the CARM offset makes us lose lock.  I think the problem is that the UGF servo demod phase is changing as we are changing offsets, filters and error signals.  We see that the I-phase is servoed successfully to 0dB, but that the Q-phase is starting to move around by 30 degrees or more.  We either need to monitor this much more closely, and add the changing demod phases to the carm_up script, or we need to go back to the sum-of-squares situation that we had last week.  Note that we saw failures with that method for a completely separate reason:  we were getting too close to the limiters, which cause the UGF servos to glitch and the outputs jump by a significant amount.  So, the issues that we were seeing last week when we had the sum-of-squares were a different thing, that we noticed and understood later.

Anyhow, nothing too exciting and glorious tonight, but progress has been made.

Also, from some Mist simulations that Q did, Diego made a sweet plot that is now posted in the control room, so we can translate arm power to CARM offset, at various MICH offsets. 

We also took some CARM loop measurements with the new filters.  We have a little more phase than we used to, which is nice.  These traces don't have the lowpass engaged, since I was trying to see how far we could get without it.  We lost lock right after the second measurement, but I think that was to do with the UGF servos.

Attachment 2: CARM_22Jan2015.pdf
CARM_22Jan2015.pdf
  10932   Thu Jan 22 22:15:30 2015 diegoUpdateSUSCentered OpLevs

[Diego, Jenne]

We centered the OpLevs for ITMX and ITMY.
 

  10931   Thu Jan 22 18:36:11 2015 diegoUpdateIOOMC Flashes

I've been looking into the data of Jan 06 and Jan 15 taken during daytime, as the night before we left the PRC aligned in order to allow the IFO to flash; the purpose is to find out if some flashes from the IFO could propagate back to the IMC and cause it to lose lock; I will show here a sample plot, all of the others are in the archive attached.

My impression is that these locklosses of the IMC are not caused by these flashes: the signals MC_[F/L] seem quite stable until the lock loss, and I don't see any correlation with what happens to REFLDC that could cause the lockloss (apart from its drop as a consequence of the lockloss itself); in addition, in most occasions I noticed that the FSS started to go crazy just before the lock loss, and that suggests me that the lockloss source is internal to the IMC.

I can't see anything strange happen to MC_TRANS either as long as the IMC is locked, no fluctuations or weird behaviour. I also plotted the MC_REFL_SUM channel. but it is too slow to be useful for this kind of "hunt".

Attachment 1: 1104612646_zoom_1.png
1104612646_zoom_1.png
Attachment 2: elog.tar.bz2
  10930   Thu Jan 22 15:35:41 2015 diegoUpdateSUSBounce/Roll Measurements

I measured the bounce/roll frequencies for all the optics, and updated the Mechanical Resonances wiki page accordingly.

I put the DTT templates I used in the /users/Templates/DTT_BounceRoll folder; I wrote a python script which takes the exported ASCII data from such templates and does all the rest; the only tricky part is to remember to export the channel data in the order "UL UR LL" for each optic; the ordering of the optics in a single template export is not important, as long as you remember it...

Anyhow, the script is documented and the only things that may need to be modified are:

  • lines 21, 22: the "starting points" FREQ_B and FREQ_R (to accomodate noisy or bad data, as ETMX was for the Roll part in both the measurements I took);
  • line 72: the parameters of the slepian window used to average the data: the first one is the most important and indicates how much averaging will happen; more averaging means less noise but broader and lower peaks, which shouldn't be a big issue since we care only about the peak position, not its amplitude; however, if the peak is already shallow, too much averaging will make things worse instead of better;
  • lines 110, 118: the initial guess for the fit parameters;

The script is in scripts/SUS/BR_freq_finder.py and in the SVN. I attach the plots I made with this method.

Attachment 1: BR_Jan2015.tar.bz2
  10929   Thu Jan 22 03:21:24 2015 JenneUpdateLSCLocks with large MICH offsets

[Jenne, Diego, EricQ]

Tonight we worked on the acquisition sequence (including re-re-re-commissioning the UGF servos, hopefully for the last time...) for the PRFPMI with large MICH offsets. 

The procedure is all in the carm_up script, as far as things work.

We had some locklosses, but they were mostly due to non-carefulness on my part during the transitions between error signals, or the UGF servos getting upset because the oscillator peaks had gotten lost in the noise.  The one that I show here is our very last one of the night, where we are hitting the rails for the MICH signal, which is then causing the other loops to have to do weird things to try to compensate, and they lose lock.

Here also is a StripTool shot during that lock stretch.  I was in the middle of increasing the MICH offset to 75% of the fringe.  The yellow trace (called MICH_B_MON) is ASDC/POPDC normalized so that it always goes 0-1.  I was pleased to see that perhaps REFL11I and AS55Q are turning over, although as Q will tell us in a more detailed elog tomorrow, having a large MICH offset does weird things and moves the DARM zero-point.  So, maybe we aren't actually anywhere awesome yet. 

After some MICH offset, the maximum arm power is always going to be about 50, so arm powers of 8 or 10 equates to 100 pm.  We didn't get there tonight while on IR signals.


The locking sequence is now something like this:

  • Lock carm and darm on ALS, find resonances, move to 3 counts (roughly 3nm) offset.
  • Set PRMI up to acquire on REFL33I and ASDC/POPDC at 25% MICH fringe.  (After a while, I assume perhaps because the alignment is no longer tip-top, I have been by-hand reducing the MICH offset from -700counts which is 25% to -200counts, and then immediately putting it back to -700 after the PRMI acquires.)
  • Engage all 4 UGF servos
  • Reduce the CARM offset a bit, to 1.0 count, which gives arm powers of about 0.4 (with 50 being the max possible)
  • Transition CARM from ALS to sqrtInvTrans
  • Transition DARM from ALS to DC trans:  (TRY-TRX)/(TRX+TRY)
  • Reduce the oscillator amplitudes of the UGF servos
  • Reduce the CARM offset to powers of about 1
  • Ramp to 50% MICH fringe

After this, we tried a few times to lower the CARM offset, but kept losing lock, I think because the UGF servos went crazy.  The final lock, shown above, we lost because the MICH output was hitting the rails.

The problem with the MICH servo right now is the low SNR of the POPDC being used to normalize ASDC.  The control output is enormous, even if we have the 400Hz lowpass on.  We need to rethink our MICH servo, starting with a lower UGF, so that we're not injecting all this sensing noise all over the place. 


For tomorrow:

  • Re-look at MICH loop, to prevent sensing noise injection.
  • How does the large MICH offset affect our zero points for CARM and DARM?  Can we stay on DC transmission signals through 30 or 100 pm?
  • What to do next?  One or two of the locklosses were because the CARM detuned double cavity pole wasn't de-Q-ed enough, so still hit 0dB and created an unstable unity gain point.  Can we go to higher MICH offset, maybe 75%? 
  • Still need to figure out where our missing phase is for our LSC loops.  CARM and DARM are short on phase, and we could definitely use some more.  So, I will work on trying to give us filters that don't eat too much phase, but we still need to find that missing ~14 deg. 
  10928   Thu Jan 22 02:56:07 2015 ericqUpdateIOOFSS input offset adjusted

Since Rana's overhaul of the IMC, the FSS input offset had been sitting at zero. 

Over the last day or so, I had noticed the MC refl wall striptool trace looking noisier, and earlier this evening, we were suffering from a fair amount of downtime due to IMC unlocks, and failure to autolock for times on the order of ten minutes. 

While we had used ezcaservo for this in the past, I set the FSS offset manually tonight. Namely, I popped open a dataviewer trace of MC_F, and scanned the FSS offset to make MC_F go to zero. It required a good amount of offset, 4.66 V according to the FSS screen. I did this while the FSS slow servo was on, which held the FSS Fast output at zero. 

That was four hours ago; MC_F is still centered on zero, and we have not had a single IMC unlock since then. The MC refl trace is thinner too, though this may be from nighttime seismic. 

  10927   Wed Jan 21 15:27:47 2015 JenneUpdateLSCFixed LSC model bug

This problem with the CARM loop last night was the fault of a bug that I had put into the LSC model last week.  When I gave the input matrix and normalization matrix double rows, I had put the goto tags for the CARM normalization matrix rows backwards.  So, even though I thought I was not normalizing CARM, in fact I was normalizing by POPDC, which was near zero since the PRM was misaligned.

Anyhow, found, fixed, currently locking, and all seems well.

  10926   Tue Jan 20 21:58:04 2015 diegoUpdateLSCSome locking, may need to modify UGF part again

 

Quote:

I have been playing with the IFO tonight.  Mostly, I wanted to make sure that all of the scripts for the carm_cm_up sequence were working, and they seem to all be fine. 

I also turned on all 4 UGF servos.  My big ah-ha moment for the night is that the excitation is multiplied by the gain multiplier.  This means that if the UGF servo is multiplying by a small number (less than 1), the excitation will get smaller, and could get small enough that it is lost in the noise.  Now the error signal for the UGF servo is very noisy, and can dip to zero.  Since we can't take the log of zero, there are limiters in the model, but they end up giving -80dB to the error point of the UGF servo.  This makes it all freak out, and often lose lock, although sometimes you just get a weird step in the UGF servo output. 

Anyhow, we need to be mindful of this, and offload the UGF servos regularly.  I think the better thing to do though will be to divide the excitation amplitude by the gain multiplier.  This will undo the fact that it is multiplied by that number, so that the number of counts that we put into the excitation amplitude box is what we expect. 

 

After some brainstorming with Jenne and Q, both the model and the medm screen have been modified: the entire block "Test1 - injection of the excitation - Test2" has been moved after the servo output. In this way we avoid completely the multiplication problem without having to perform divisions that could lead to division-by-zero problems. Because of how the logic is done now, one UnitDelay block had to be inserted before each one of Test1 and Test2.

 

Since the UGF Servo has been heavily modified lately, I'll post the current status of the model (as an attachment, as inpage images lose too much quality).

 

Attachment 1: UGF_SERVO_MDL.png
UGF_SERVO_MDL.png
  10925   Tue Jan 20 20:03:17 2015 JenneUpdateLSCALS lock not staying?

The Xarm ALS has been a little funky today. 

First, the green and the arm-axis would not stay co-aligned.  I'm not sure which was moving (although neither ITMX nor ETMX seemed to be moving very much according to their oplevs and OSEMs).  I went to the Xend table and jiggled the mounts for the steering optics, in case one was loose or something.  None were.  However, the transmission quit jumping around by a factor of 2 after that.   The beatnote alignment on the PSL table was also bad, so I tweaked that alignment up for the Xarm.  There were some not connected cables and fibers blocking the access to the X beatnote area, so those are up on top of the PSL table.

Anyhow, I haven't locked the individual arms, but I cannot hold the lock with CARM/DARM.  The CARM output keeps hitting the threshold for the locking watchdog, which turns off the lock.  Obviously I could just increase this threshold, but the right thing to do is figure out why the Xarm ALS is so noisy today, and why it wants to output such a large control signal to maintain the lock.

  10924   Tue Jan 20 19:32:06 2015 JenneUpdateSUSSUS Drift restore scripts

I made little scripts to go with the sus driftmon buttons, that will servo the alignment sliders until the susyaw and suspit values match the references on the driftmon screen. 

  10923   Tue Jan 20 15:09:01 2015 JenneUpdateLSCLSC model change implemented

 

Quote:

Brain not working anymore now that it's ~4am, but I need to rethink and recheck to make sure that the PD whitening triggering is still okay and working.  Or maybe we can remove it, and just include that in the scripts, as Rana has been suggesting for ages.  Thoughts for tomorrow.

LSC whitening triggering was not working, because of the implementation of the double-rows for the input matrix.  I have modified the c-code that looks at the input matrix and triggers, and decides when to turn on the PD whitening, so that it now works.

  10922   Tue Jan 20 09:32:19 2015 SteveUpdatePEMdusty Monday morning

Yesterday morning was dusty. I wonder why?

The PRM sus damping was restored this morning.

Attachment 1: dustyMorning.png
dustyMorning.png
  10921   Tue Jan 20 02:39:49 2015 ericqUpdateASCQPD ASC saga continues.

Although the QPD loops are less than ideal right now, I made changes to the ASC model to trigger the QPD loops on and off politely, depending on TRX and TRY. The settings are exposed on the ASC screen. However, I have not yet exposed the FM triggering that I also set up to make sure the integrator doesn't misbehave if the arm loses lock. We probably don't want to trigger them on at anything lower than arm powers of about 1.0. 

I've tested the triggering by randomly turning LSC mode on and off, and making sure that the optics don't recieve much of a kick as the QPD loops engage a few seconds after the LSC boosts do, or when lock is lost. This works as long as there isn't much of a DC offset befire the loops are engaged. (Under 20 counts or so is fine)

As a side note, I was going to use the TRIG_SIG signals sent via the LSC model via SHMEM blocks for the ASC triggering, but oddly, the data streams that made it over were actually the MICH and SRCL TRIG_SIGs, instead of XARM and YARM as labelled. I double checked the simulink diagrams; everything seemed fine to me. In any case, ASS was already recieving TRX and TRY directly via RFM, so I just piped those over to the ASC block. This way is probably better anyways, because it directly references the arm powers, instead of the less obvious LSC triggering matrix. 

  10920   Mon Jan 19 18:27:16 2015 ericqUpdateASCQPD ASC saga continues.

Herein, I will try to paint a more thorough picture of this whole QPD ASC mess. 


Motivation for QPD ASC loops:

  • We would like to use the QPDs as a DC arm pointing reference during locking attempts, or over multiple days, if possible. 
  • It would be nice if the QPDs could complement the oplevs to reduce angular motion of the cavity. 
  • We must not make the single arm longnitudinal noise or RIN worse, because anything observable in the single arm case will be catastrophic at full sensitivity

Actuation design:

As mentioned in a previous ELOG, in single arm lock, I measured the QPD response with respect to the calibrated oplev signals. They were:

YARM
ETMY: QPD PIT / OPLEV PIT =   22.0 count/urad
      QPD YAW / OPLEV YAW =   17.1 count/urad
ITMY: QPD PIT / OPLEV PIT =   -6.0 count/urad
      QPD YAW / OPLEV YAW =    5.9 count/urad
XARM
ETMX: QPD PIT / OPLEV PIT = 16.6 count/urad
      QPD YAW / OPLEV YAW = -9.3 count/urad
ITMX: QPD PIT / OPLEV PIT =  4.0 count/urad
      QPD YAW / OPLEV YAW = -6.0 count/urad

For reference, one microradian of either ITM or ETM motion produces about 60um of ETM beam spot displacement, compared to the spot size of ~5mm. 

However, given the lenses on the end tables that are used for green mode matching, that the IR transmitted beam also passes through, the QPDs are not directly imaging the ETM spot position; if they were, they would have equal sensitivity to ITM and ETM motion due to our flat/curved arm geometry. 

From this data, I calculated the actuation coefficients for each DoF as A_{ETM} = \frac{d_{ETM}}{\sqrt {d_{ETM}^2 + d_{ITM}^2}}, and similarly for the ITMs, where the d's come from the table above. However, it occurs to me that maybe this isn't the way to go... I'll mention this later. 


Loop design:

Up until now, at every turn, I had not properly been thinking about how the oplev loop plays into all of this. I went to the foton filters, and grabbed the loop and plant models for the ETMY oplev, and constructed the closed loop gain, 1/1+G, and the modified plant, P/1+G, which is what the ASC loop sees as its plant. 

Here, the purple trace explains all of the features I was confused about earlier. 

With this in hand, I set up to design a loop to satisfy our motivations. 

  • Bounce/roll mode notches to avoid exciting them
  • 1/f UGF crossing at a few Hz, limited by the gain margin at ~10Hz, which is where the phase will hit 180, due to the notches
  • 4th order Elliptic lowpass at 100, to avoid contaminating the longnitudinal signals
  • 1/f^2 at low frequencies for DC gain

To do this, I inverted the oplev closed loop plant pole around 4Hz to smooth the whole thing out. Here's a comparison of the measured OLG with what I modelled. 

There's a little bit of phase discrepency around 10Hz, but I think it looks about right overall. 


Evaluation:

So, here's the part that counts: How does this actually perform? I took spectra of the QPD error signals, the relevant OpLev signals as out of loop sensors, the PDH error signal and transmitted RIN while single arm locked, with this loop off, and on for all 4 DoFs simultaneously. 

Verdict:

  • In-loop signals get small, unsurprisingly.
  • Cavity signals unchanged. yes
  • ITM oplev signals are unchanged (and not plotted, to not clutter the plots (This isn't surprising since the loops mostly actuate on the ETMs).
  • ETM oplev signals get smaller around the 3Hz peak, but are louder in other bands.no 

This is what makes me think I may need to revisit the actuation matrix. If I did it wrong, I am driving the "invisible" quadrant of the cavity angular DoFs, and this could be what is injecting noise into the oplevs. 


Conclusion:

In the end, I have a better understanding of what is going on, and I don't think we're quite there yet.  

Attachment 1: oplevPlant.pdf
oplevPlant.pdf
Attachment 2: loopDesignComparison.pdf
loopDesignComparison.pdf
  10919   Mon Jan 19 11:03:32 2015 manasaUpdateGeneralRF connections to adder

RF connections to the beat note adder on the IOO rack have been undone and it should be good to proceed with any locking attempts.

Quote:

I was around the PSL table and the X end table today.

X end table:

I was at the X end table making some distance measurements for the telescope to the fiber coupler. I have NOT moved anything as yet.

PSL table:

I wanted to get some data to look at the beat note frequency discrepancy between the green and IR. 

I tried making measurements using the spectrum analyzer at the PSL table but the MC was getting unlocked quite often with PSL enclosure open and HEPA on high. 

So I switched off the power supplies to the RFPDs (3 of them) on the PSL table and disconnected the Xmon input to the adder (at the IOO rack) which brings the green beat note signals to the conrol room. I connected the fiber RFPD output to the adder and took a bunch of measurements of the green and IR beat notes from the spectrum analyzer in the control room. I have NOT undone this setup assuming there are no locking plans for tonight and I will come back tomorrow. 

If anyone is planning to do some locking in the meantime, you can undo the connections keeping in mind to turn off the power to the RFPDs before you do so.

P.S. I walked in this morning to PSL FSS slow actuator at 0.89 and brought it down to zero before making measurements.
I also touched the steering mirrors that are part of the Y green beat note alignment while looking at the amplitude of the RF signal on the spectrum analyzer.

 

  10918   Sat Jan 17 15:07:28 2015 manasaUpdateGeneralBeat note frequency discrepancy

I was around the PSL table and the X end table today.

X end table:

I was at the X end table making some distance measurements for the telescope to the fiber coupler. I have NOT moved anything as yet.

PSL table:

I wanted to get some data to look at the beat note frequency discrepancy between the green and IR. 

I tried making measurements using the spectrum analyzer at the PSL table but the MC was getting unlocked quite often with PSL enclosure open and HEPA on high. 

So I switched off the power supplies to the RFPDs (3 of them) on the PSL table and disconnected the Xmon input to the adder (at the IOO rack) which brings the green beat note signals to the conrol room. I connected the fiber RFPD output to the adder and took a bunch of measurements of the green and IR beat notes from the spectrum analyzer in the control room. I have NOT undone this setup assuming there are no locking plans for tonight and I will come back tomorrow. 

If anyone is planning to do some locking in the meantime, you can undo the connections keeping in mind to turn off the power to the RFPDs before you do so.

P.S. I walked in this morning to PSL FSS slow actuator at 0.89 and brought it down to zero before making measurements.
I also touched the steering mirrors that are part of the Y green beat note alignment while looking at the amplitude of the RF signal on the spectrum analyzer.

  10917   Sat Jan 17 01:10:36 2015 JenneUpdateLSCSome locking, may need to modify UGF part again

I have been playing with the IFO tonight.  Mostly, I wanted to make sure that all of the scripts for the carm_cm_up sequence were working, and they seem to all be fine. 

I also turned on all 4 UGF servos.  My big ah-ha moment for the night is that the excitation is multiplied by the gain multiplier.  This means that if the UGF servo is multiplying by a small number (less than 1), the excitation will get smaller, and could get small enough that it is lost in the noise.  Now the error signal for the UGF servo is very noisy, and can dip to zero.  Since we can't take the log of zero, there are limiters in the model, but they end up giving -80dB to the error point of the UGF servo.  This makes it all freak out, and often lose lock, although sometimes you just get a weird step in the UGF servo output. 

Anyhow, we need to be mindful of this, and offload the UGF servos regularly.  I think the better thing to do though will be to divide the excitation amplitude by the gain multiplier.  This will undo the fact that it is multiplied by that number, so that the number of counts that we put into the excitation amplitude box is what we expect. 

  10916   Fri Jan 16 20:37:52 2015 diegoUpdateLSCUGF servo now linear again

I found an error in the model of the UGF servos, I have now corrected it; for future reference, now the division between TEST2 and TEST1 is properly done with complex math: given

TEST1 = a + i b\hspace{.5cm},\hspace{.5cm}}TEST2 = c + id

we have that TEST3:

 

TEST3 = \frac{TEST2}{TEST1} = \left(\frac{ac+bd}{a^2+b^2}\right) + \left(\frac{ad-bc}{a^2+b^2} \right)i

 

TEST3 is the actual signal that is now phase rotated to select only the I signal while rejecting the Q one.

 

All the updates to the model, the screens and the script have been SVNed.

  10915   Fri Jan 16 20:01:32 2015 JenneUpdateLSCLSC model change implemented

Nope, I used the script.

Yesterday's changes were mostly to the generateLSCscreen/C1LSC_OVERVIEW_INPUT_MATRIX.adl sub-screen.  The UGF servos were added earlier in the week to the LSC screen in the generateLSCscreen/C1LSC_OVERVIEW_SERVOS.adl sub-screen.

  10914   Fri Jan 16 18:46:15 2015 KojiUpdateLSCLSC model change implemented

Was the screen modified directly on LSC_OVERVIEW.adl?
Even if so, that's OK. I'll incorporate the change to the screen making script once I'm back.

  10913   Fri Jan 16 18:09:09 2015 JenneUpdatePSLPMC autolocker not running?

Have we been running the PMC autolocker lately?  I can't remember, and I also can't find where it might be running.  It's not on megatron, either in the crontab or Q's new /etc/init place.  It's also not on op340m. 

Anyhow, what prompted this was that the PMC transmission has been incredibly fuzzy today.  On the StripTool it looks like it was fine until about -7 hours ago, when it lost lock.  Then Diego relocked it around -3 hours ago, and it's been fuzzy ever since. It was unlocked again for about 15 minutes about 45 minutes ago, and when I relocked it, it was even more fuzzy.

The FSS slow is almost exactly zero, the PMC's PZT is not at the edge of the range, the FSS PC drive RMS has been both high and low, and the PMC fuzz level has just been consistently high.  I was checking the parameters in the PMC phase shifter screen, and looked up the autolocker to see what the nominial values are supposed to be. 

For the RFADJ value, the autolocker sets it to 2.0, and after it locks increases it up to 4.5.  I found the value at 2.0, and the autolocker isn't running, which made me wonder if the autolocker died sometime after it set the value low, but before it could detect lock and reset the value to high.  (Actually, after lock it sets the value to whatever is in the channel C1:PSL-STAT_PMC_NOM_RF_ADJ, which is 4.5).

Anyhow, I manually set the RFADJ value to 4.5, and the PMC transmission immediately improved. 

 

EDIT, 8pm, JCD:  Rana reminded me that he attached a screenshot back on 30Dec2014 (http://nodus.ligo.caltech.edu:8080/40m/10849), which I had ignored earlier because the parameters weren't written in text.  My bad.  Anyhow, after the New Year's tune-up, the RFADJ should be 6.0.  I have set it so, and also changed the C1:PSL-STAT_PMC_NOM_RF_ADJ chan to be 6.0.

Attachment 1: PMCfuzz.pdf
PMCfuzz.pdf
  10912   Fri Jan 16 04:14:38 2015 ericqUpdateCDSUnexpected CDS behavior

EDIT: Sleepy Eric doesn't understand loops. The conditions for this observation included active oplev loops. Thus, obviously, looking at the in-loop signal after the ASC signl joins the oplev signal will produce this kind of behavior. 


After some talking with Rana, I set out on making an even better-er QPD loop. I made some progress on this, but a new mystery halted my progress. 

I sought to have a more physical undertanding of the plant TF I had measured. Earlier, I had assumed that the 4Hz plant features I had measured for the QPD loops were coming from the oplev-modified pendulum response, but this isn't actually consistent with the loop algebra of the oplev servos. I had seen this feature in both the oplev and qpd error signals when pushing an excitation from the ASC-XARM_PIT (and so forth) FMs. 

However, when exciting via the SUS-ETMX-OLPIT FMs (and so forth), this feature would not appear in either the QPD or oplev error signals. That's weird. The outputs of these two FMs should just be summed, right before the coil matrix. 

I started looking at the TF from ASC-YARM_PIT_OUT to SUS-ETMY_TO_COIL_1_2, which should be a purely digital signal routing of unity, and saw it exhibit the phase shape at 4Hz that I had seen in earlier measurements. Here it is:

I am very puzzled by all of this. indecision Needs more investigation.

Attachment 1: digitalProblem.pdf
digitalProblem.pdf
  10911   Fri Jan 16 04:14:05 2015 diegoUpdateLSCUGF servo now linear again

UGF Servos' commissioning still going on, updates of today:

  • on Rana's suggestion, we don't use anymore the Q-signal rejection at the level of Phase 1 and Phase 2; instead, a proper complex division is made between those two signals (with a check in case of zero); then the resulting signal is demodulated with a new Phase 3, which is the one used to select the I signal while zeroing the Q one; changes to the model and the screens have been made;
  • a new evaluation of all the parameters for the four servos has been made; aside for the new phase, and the zeroing of the other phases (because they are not used anymore for the selection), the parameters are not dissimilar from the prevoius ones
  • the PRCL and MICH servos seem sufficiently stable;
  • CARM and DARM are stable only for a short amount of time; what usually happens is that one of the two starts drifting in one random direction, and usually the other one follows shortly after; it is not clear if there is a relation or if they stop being stable after a similar amount of time; I still noticed a few lowest limits appearing in the input signal, which should be avoided; I'll check the model again tomorrow;
  • the weird thing about CARM and DARM is that at the same time when one of them starts drifting, its I and Q signals begin to be comparable; when the servo is shut off, they resume their normal state;
  • an increase in the excitation gain improves the separation of I and Q and also reduces their variations, but a high peak in the loop due to this might not be a good idea.

 

  10910   Fri Jan 16 03:31:35 2015 JenneUpdateLSCLSC model change implemented

Okay, it has taken me almost exactly 12 hours (with a dinner break), but I have implemented this change.

Everything was svn-ed before I did things, and then again afterward.

Here is the "before" screenshot of the LSC model:

And here is afterward:

If you look extra carefully, you will see that it matches my proposal from http://nodus.ligo.caltech.edu:8080/40m/10904 .  I have made the change for DARM, MICH, PRCL, SRCL and CARM.  I did not alter XARM, YARM or MC.  Also, the CESAR stuff was taken out of the CARM area, since this is now a more generalized version of the same thing.


I have also checked and modified all of the scripts that I could think of, as well as all of the ifoconfig burt .req and .snap files that I could think of.  I also ran the carm_cm_up.sh script once, and it seems to work fine.  All of the transition scripts that are listed below (which are the only ones used currently in the sequence) now use the new error signal blending scheme.

Scripts:

  • Lock_ALS_CARMandDARM.py
  • ALSfindIRresonance.py
  • ALSwatch.py
  • carm_cm_down.sh
  • carm_cm_up.sh
  • CheckPRMIlock.py
  • Transition_MICH_REFL33Q_to_ASDC.py
  • Transition_CARM_ALS_to_TransInvSqrt.py
  • Transition_DARM_ALS_to_DCtrans.py
  • UGFup.py
  • UGFdown.py

Burts (listed are the .req files, but I also checked the .snap files, hand-editing the matrix element numbers where needed if I wasn't in the right config to do a save):

  • C1configure_Yarm.req
  • C1configure_YarmALS.req
  • C1configure_Xarm.req
  • C1configure_XarmALS.req
  • C1configure_CARM.req
  • C1configure_DARM.req
  • C1configure_PRM_forCARMdarm.req
  • C1configure_PRM_SBres.req
  • C1configure_PRM_Carr.req
  • C1configure_PRY.req
  • C1configure_DRM.req
  • C1configure_SRM.req

I also modified the screens for the input matrix and for the normalization matrix.  I created a new screen for the final blending matrices (which are all 2x1's), and I also modified the LSC_OVERVIEW screen. 

The input matrix and normalization matrix screens have colored bars that tell you whether a row is in use or not.  If the background to the row is the blue of the whole screen, that row is not being used.

The LSC screen has new hand-modified Kissel Buttons.  I wanted to show the total PD error signal that is being used, regardless of what row (A or B) it is on at that time.  So, I have collapsed the rows so that DARM_A and DARM_B are overlapped, even though they are actually rows 1 and 2 of the matrix.  The PD should only show up green on the LSC screen if that row is in use (so, if you are prepping a row, but aren't using it yet, you won't see those elements in the matrix).  Anyhow, the point is that the LSC overview part of things shouldn't look any different than before.


Brain not working anymore now that it's ~4am, but I need to rethink and recheck to make sure that the PD whitening triggering is still okay and working.  Or maybe we can remove it, and just include that in the scripts, as Rana has been suggesting for ages.  Thoughts for tomorrow.

  10909   Thu Jan 15 19:01:30 2015 ericqUpdatePSLPMC realigned

PMC realigned again... The transmission was down to 0.70, and the MC was having a hard time trying to autolock.

  10908   Thu Jan 15 18:57:41 2015 ericqUpdateASCTransmon QPD loops live

I've measured the sensing for each of the arms, by using our calibrated oplevs, in terms of QPD counts per micron. It is:

YARM
ETMY: QPD PIT / OPLEV PIT =   22.0 count/urad
      QPD YAW / OPLEV YAW =   17.1 count/urad
ITMY: QPD PIT / OPLEV PIT =   -6.0 count/urad
      QPD YAW / OPLEV YAW =    5.9 count/urad
XARM
ETMX: QPD PIT / OPLEV PIT =   16.6 count/urad
      QPD YAW / OPLEV YAW =   -9.3 count/urad
ITMX: QPD PIT / OPLEV PIT =    4.0 count/urad
      QPD YAW / OPLEV YAW =   -6.0 count/urad

In the absence of a lens, the QPD would be significantly more sensitive to cavity axis translation than tilt, and thus about equally sensitive to ITM and ETM angle. However, there are lenses on the end tables. I didn't go out and look at them, but found some elogs from Annalisa that mentioned 1m focal length lenses. Back-of-the-envelope calculations convince me that this can plausibly lead to the above sensitivity ratios.

I used these quantities to come up with an actuation matrix for the ASC loops, and measured the effective plant seen by the FM, fitted it to some poles( looks like zpk([],-2*pi*[1.47+3.67i,1.47-3.67i],160); ), and designed a control servo. Here is the designed loop:

The servo works on both arms, both DoFs. A DTT measurement agrees with the designed loop shape, up to a few degrees, which are probably due to the CDS delay. The RMS of the QPD error signals goes down by about 20dB, and are currently dominated by the bounce mode, so maybe we can try to sneak in some resonant gain...?

Once we confirm that they work when locking, we can write up and down lines into the locking scripts...

Attachment 1: loopDesign.pdf
loopDesign.pdf
  10907   Thu Jan 15 18:30:18 2015 JenneUpdateComputer Scripts / ProgramsInstalled kerberos on Rossa

 

Quote:

WARNING: since the workstations are all shared user, if you forget to kdestroy the next user can commit under your user ID.  It might be good to set the timeout to be something much shorter than 24 hours, like maybe 1, or 2.

Good call.  I added a line ticket_lifetime = 3600, which should make it destroy the credentials after an hour.

  10906   Thu Jan 15 18:10:19 2015 jamieUpdateComputer Scripts / ProgramsInstalled kerberos on Rossa
Quote:

I have installed kerberos on Rossa, so that I don't have to type my name and password every time I do an svn checkin, since I'm making some modifications and want to be sure that everything is checked in before and afterwards. 

I ran sudo apt-get install krb5-user.  I didn't put in a default_realm when it prompted me to during install, so I went into the /etc/krb5.conf file and changed the default_realm line to read default_realm = LIGO.ORG

Now we can use kinit, but we must (as usual) remember to kdestroy our credentials when we're done.

As a reminder, to use:

> kinit albert.einstein

Password for albert.einstein@LIGO.ORG: (type your pw here)

When you're finished, run

> kdestroy

The end.

WARNING: since the workstations are all shared user, if you forget to kdestroy the next user can commit under your user ID.  It might be good to set the timeout to be something much shorter than 24 hours, like maybe 1, or 2.

  10905   Thu Jan 15 18:06:34 2015 JenneUpdateComputer Scripts / ProgramsInstalled kerberos on Rossa

I have installed kerberos on Rossa, so that I don't have to type my name and password every time I do an svn checkin, since I'm making some modifications and want to be sure that everything is checked in before and afterwards. 

I ran sudo apt-get install krb5-user.  I didn't put in a default_realm when it prompted me to during install, so I went into the /etc/krb5.conf file and changed the default_realm line to read default_realm = LIGO.ORG

Now we can use kinit, but we must (as usual) remember to kdestroy our credentials when we're done.

As a reminder, to use:

> kinit albert.einstein

Password for albert.einstein@LIGO.ORG: (type your pw here)

When you're finished, run

> kdestroy

The end.

  10904   Thu Jan 15 14:28:14 2015 JenneUpdateLSCLSC model change idea

Something that kind of drives me crazy with our current LSC model setup is that I can't make "finished" error signals before blending them.  The blending happens before the normalization matrix, and there is no place to put an offset to help match a new error signal to the current offset.  So.  While I'm sure this is not going to be immediately popular, here's a cartoon of a proposed model change to the LSC. 

The most important difference between this and the ramping matrix that is used at the sites is that you can put in offsets before the blend.  Also useful is the fact that the normalization can happen before the blend.  This proposal would make the LSC input matrix  and the normalization matrix have twice as many rows, and add an extra "selector matrix" just before the triggering at the error point of the loops. 

I've only drawn one degree of freedom in my cartoon, but assume that they all have the same capability (maybe we don't have to do XARM, YARM and MC this way).  One row is currently being used for the error signal, while the other row is just used to prep a new singal.  For a first transition (say, ALS to DC transmission), maybe the ALS signals are on row 1, and the DC trans is on row 2.  Once the transition is complete, row 1 is available to prep for the next transition (such as AS55Q).

 

Thoughts?  Is there a better way to achieve what I'm going for here?

Attachment 1: SwitchableErrorSignalProposal_Jan2015.pdf
SwitchableErrorSignalProposal_Jan2015.pdf
  10903   Thu Jan 15 04:41:01 2015 JenneUpdateLSCThoughts on going forward with variable finesse

[Jenne, Diego]

Life would be easier with the UGF servos working.  As Diego already elogged, we aren't sure why the demod phases are changing, but that is certainly causing the I-signals to dip below zero, which the log function can't handle (there is a limiter before the log, so that the signal can't go below 1e-3).  Anyhow, this is causing the UGF servos to freak out, so we have not been using them for tonight's locking.

Our goal tonight was to see if we could introduce a nice big MICH offset, and then lower the CARM offset while keeping the arms locked on ALS.  We hope to see some kind of sign of a PDH signal in some RF PD. 

In the end, the highest we got to was -460 MICH offset counts, which we think is about 29nm (if our rough calibration is accurate). The MICH half fringe should be 188nm. With this offset, we got down to 0.3 CARM offset counts while locked on ALS.  We think that this is around 300pm, plus or minus a lot.  Note that while yesterday I had a pretty easy time getting to -660 counts of MICH offset, tonight I struggled to get past -200.  The only way we ended up getting farther was by lowering the CARM offset.  Although, as I type this, I realized that last night's work already had a lower CARM offset, so maybe that's key to being able to increase the MICH offset. 

We watched REFL11I and REFL11I/(TRX+TRY) on striptool, but we didn't see any evidence of a PDH signal.  We lost lock when I tried to transition CARM over to REFLDC, but I wasn't careful about my offset-setting, so I am not convinced that REFLDC is hopeless.

So.  Tonight, we didn't make any major locking progress (the MC started being fussy for about an hour, right after I ran the LSC offsets script, just before we started locking in earnest).  However, we have some ideas from talking with Rana about directions to go:

* Can we transition CARM over to REFL11I, and then engage the AO path?

* Then, while the MICH offset is still large, can we transition DARM over to POX or POY, actuating on a single arm?  If CARM is totally suppressed, this is DARM-y.  If CARM doesn't have the AO path yet, this is halfsy-halfsy, but maybe we don't care.

* Then, can we lower the MICH offset and transition back to a REFLQ signal?

* Separately, it seemed like we kept losing PRC lock due to PRC motion.  If the MICH offset is very large, are we sideband-limited at the POP port, such that we can use the POP DC QPD?  Is it even worth it?

 

MICH calibration:

A single mirror (ITM) moving by lambda/2, in the MICH-only situation is the full range, from bright to dark fringe.  So, half fringe should be lambda/4, or about 133nm.  If we are thinking about pushing on the BS, there's an extra factor of sqrt(2), so I think the half fringe should be at 188nm of BS motion.

When we had MICH locked on ASDC/POPDC, we put in a line at 143.125Hz, at 20 counts to (0.5*BS-0.2625*PRM), so a total of 10 counts to the BS at 143Hz.  Given the BS calibration in http://nodus.ligo.caltech.edu:8080/40m/8242, this is 10.1pm of actuation.  We saw a line in the error signal of 0.1 counts, so we infer that the MICH error signal of ASDC/POPDC has a calibration of 94pm/count. This number was invariant over a few different MICH offsets, although the ones I measured at were all below about 100 counts of MICH offset, so maybe this number changes as we start to get farther from the MICH dark fringe.

 

IFO left flashing (all mirrors aligned except SRM) in case anyone wants fresh data for that.

  10902   Thu Jan 15 03:18:11 2015 diegoUpdateLSCUGF servo now linear again

The UGF servos were recommissioned today:

  • suitable values of frequency, excitation, phases and gain were found;
  • the phases were chosen in order to maximize the I signal and suppress the Q one;
  • the servos seemed sufficiently stable when in a quiet state, but they didn't performed well in other cases;
  • I also found out that DARM & CARM and MICH & PRCL are maybe too much coupled, but that could be actually due to the main loops rather than the UGF ones;
  • however, after some weird rampings with no apparent reasons, and after some quite bad and glitchy step responses, I found out that the effect of the chosen phases vanished: the I and Q signals were of the same order of magnitude again, probably causing the bad performance;
  • Jenne and I tried to increase che SINGAINs and COSGAINs (but keeping them equal to each other): this has the good effect of separating more the I and Q signals, but it's just a zoom effect: there still are mixing effects and, more important, some zero-crossings into negative values that cause the signal going into the servo to go crazy.

Our idea is that we need with some thinking about these servos and most of all try to figure out all this phase thing before we can start to trust the servos to be used for locking.

  10901   Wed Jan 14 19:27:09 2015 ericqUpdateLSCUGF servo now linear again

The UGF servos have been moved to the control point, are are once again totally linear!

  10900   Wed Jan 14 03:42:31 2015 JenneUpdateLSCThoughts on going forward with variable finesse

[Jenne, Rana]

We tried locking with the variable finesse MICH offset technique again today. 

A daytime task tomorrow will be to figure out where we are in MICH and CARM offset spaces.  This will require some thinking, and perhaps some modelling.

We were using the UGF servos and checking out their step resonses, and had the realization that we don't want the gain multiplication to happen before the offsets are applied, in the case of MICH and CARM.  Otherwise, as the UGF servo adjusts the gain, the offset is changed.  I think this is what ChrisW and I saw earlier on in the evening, when it seemed like the CARM offset spontaneously zoomed toward zero even though I didn't think I was touching any buttons or parameters.  Anyhow, we no longer used the MICH and CARM UGF servos for the rest of the night.  We need to think about where we want the offset to happen, and where we want the UGF servo multiplication to happen (maybe at the control point, with a very low bandwidth?) such that this is not an issue.

Also, I'm no longer sure that the sqrt(I^2 + Q^2) instead of the usual demodulation is going to work for the UGF servos (Q made this change the other day, after we had talked about it).  When the numbers going into the I and Q servo banks are small (around 1e-5), the total UGF servo gets the answer wrong by a factor of 10 or so.  If I made the "sin gain" and "cos gain" 1000 instead of the usual 1, the numbers were of the order 1e-2, and the servo worked like normal.  So, I think we were perhaps running into some kind of numerical error somehow.  We first noticed this when we lowered the DARM excitation by a factor of 10, and the servo no longer functioned.  We should take out this non-linear math and go back to linear math tomorrow.

During the evening tomorrow, we should try locking the PRMI with a large MICH offset, and then leaving CARM and DARM on ALS, and seeing how far we can get.  Is it possible to just jump over to RF signals, since we won't have to worry about the detuned cavity pole?

Tonight, the locking procedure was the same as usual, but stopping the carm_up script before it starts to lower the CARM offset at all.  Only difference was that MICH triggered FMs were 2,3,7 rather than the usual 2,6,8. 

So, assuming you have the IFO with CARM and DARM on ALS held at +3 CARM offset counts (which we think is about 3nm), and the PRMI is locked on REFL33I&Q with no offsets, here's what we did:

  • PRCL UGF servo on
  • MICH offset goes to -20
  • MICH transitions to ASDC (0.27*ASDC, then normalize by POPDC)
  • DARM UGF servo on
  • CARM offset to 1 (arms about 0.25)
  • CARM transition to SqrtInvTrans
  • Lower CARM gain to 4
  • CARM offset to 0.6 (arms about 1)
  • DARM transition to DC transmission
  • Increase MICH offset to about -650 or -670
  • Lower CARM offset, see what happens

Something else to think about:  Should we normalize our DC transmission signals by POPDC, so that the arm powers will change when we change the MICH offset (for a constant CARM offset)?

The best we got was holding for a few minutes at arm powers of 7.5, but since the MICH offset was large and the power recycling was low, this was perhaps pretty far.  This is why we need some calibration action.

Also, earlier today I copied the CARM and DARM "slide" filter module screens so that we have the same thing for MICH.  Now all 3 of these degrees of freedom have slider versions of the filter module screens, which are called from the ctrl_compact screen.

  10899   Wed Jan 14 02:11:07 2015 ranaSummaryTreasure2-loop Algebra Loopology

I show here the matrix formalism to calculate analytically the loop TF relationships for the IMC w/ both FSS actuators so that it would be easier to interperet the results.

The attached PDF shows the Mathematica notebook and the associated block diagram.

In the notebook, I have written the single hop connection gains into the K matrix. P is the optical plant, C is the Common electronic gain, F is the 'fast' NPRO PZT path, and M is the phase Modulator.

G is the closed loop gain matrix. The notation is similar to matlab SS systems; the first index is the row and the second index is the column. If you want to find the TF from node 2 to node 3, you would ask for G[[3,2]].

As examples, I've shown how to get the FAST gain TF that I recently made with the Koji filter box as well as the usual OLG measurement that we make from the MC servo board front panel.

Attachment 1: FSSloop.pdf
FSSloop.pdf FSSloop.pdf
Attachment 2: FSSloop.png
FSSloop.png
  10898   Tue Jan 13 23:17:57 2015 ChrisFrogsComputer Scripts / Programsmedm time machine

After recompiling medm with a patch for dumping screens (attached), I added a time machine to the right-click Execute menu.  It's installed under /cvs/cds/caltech/users/wipf/src/medm_time_machine. Dependencies include the python CA server module (pcaspy) and the latest nds2-client 0.11.2.  These were also installed under my users directory, to avoid interfering with other tools.

 

Attachment 1: tm.png
tm.png
Attachment 2: dump_medm_screen.patch
--- /ligo/apps/epics-3.14.12_long/extensions/src/medm/medm/utils.c.orig	2015-01-13 18:56:44.867720104 -0800
+++ /ligo/apps/epics-3.14.12_long/extensions/src/medm/medm/utils.c	2015-01-13 22:49:56.636820963 -0800
@@ -4156,6 +4156,37 @@
 	timeOffset = time900101 - time700101;
 }
 
+#if((2*MAX_TRACES)+2) > MAX_PENS
+#define MAX_COUNT 2*MAX_TRACES+2
+#else
+#define MAX_COUNT MAX_PENS
... 75 more lines ...
  10897   Tue Jan 13 18:47:20 2015 ChrisConfigurationComputer Scripts / Programsinstafoton setup

To use instafoton, right click an MEDM screen, open the Execute menu, and choose "Foton".  Then click on the EPICS channel of a filter module as displayed on the screen.

Here's how it was set up:

  • Install instafoton.py in /opt/rtcds/caltech/c1/scripts; edit paths to localize for the 40m
  • Add instafoton to the MEDM_EXEC_LIST environment variable, newly defined in /ligo/cdscfg/workstationrc.sh:
    export MEDM_EXEC_LIST="Edit this screen;medm &A &:Probe;probe &P &:Foton (Pick filter PV);/opt/rtcds/caltech/c1/scripts/instafoton.py &P &"
  10896   Tue Jan 13 15:11:30 2015 diegoUpdateLSCTransitioned to ASDC MICH (PRMI and PRFPMI)

These are the parameters of the UGF servos we used last night:

DOF / Parameters Exc. Frequency (Hz) Exc. Gain Loop Gain
DARM 110.1 0.025 -0.03
MICH n/a n/a n/a
PRCL 150.001 2.0 -0.03
CARM 115.1 0.01 -0.03

 

Some tweaking of such parameters and the commissioning of the MICH servo will be done soon; an elog post about the UGF scripts/medm screens also will be done soon.

  10895   Tue Jan 13 14:01:20 2015 manasaUpdateGeneralAUX Y + PSL beat note at 1064nm: needs work
Quote:

I'm super excited about this new frequency readback, but I'm not sure that it's reliable yet.  Without touching any settings, the readback is currently saying 78.6MHz, and is changing slightly (as is the FSS Slow Temp), so something is working.  However, the beatnote as measured on the spectrum analyzer is 158.2MHz.  So, either the calibration or the tracking or something isn't quite finished being tuned yet.

It's going to be super awesome when we have this though!!

As Jenne pointed out, this is still not fully tuned. 

For example, I found today that the frequency counter requires more power at the input >= -20dBm to measure frequency inputs < 40 MHz. Since the RFPD gives ~ -40dBm at its output, the ~20dB gain amplifier will not be enough to measure low frequencies in case the beat power at the PD drops (which is very much possible when the alignment drifts or things move around on the PSL table).  So I am shopping for an RF amplifier with higher gain to replace the current one. In the meantime, I will test the PID loop for set point frequency > 40MHz.

I have also observed the frequency difference between the measured frequencies on FC and the spectrum analyzer. I am not sure where this comes from as yet. 

At this point, the FC readout is only reliable enough to find a beatnote that is lost on the spectrum analyzer.

  10894   Tue Jan 13 13:22:54 2015 JenneUpdateGeneralAUX Y + PSL beat note at 1064nm: needs work

I'm super excited about this new frequency readbacklaugh, but I'm not sure that it's reliable yetfrown.  Without touching any settings, the readback is currently saying 78.6MHz, and is changing slightly (as is the FSS Slow Temp), so something is working.  However, the beatnote as measured on the spectrum analyzer is 158.2MHz.  So, either the calibration or the tracking or something isn't quite finished being tuned yet.cheeky

It's going to be super awesome when we have this though!!cool

  10893   Tue Jan 13 10:28:02 2015 manasaUpdateGeneralAUX Y + PSL beat note at 1064nm
Quote:
 

I plan to use the Minicircuits ZFL-1000ln that is on the IOO rack but not being used (This was used for green beatnote amplification but is not required/used anymore) to amplify the RF signal to the frequency counter.
If anyone has any objections to using this amplifier for the frequency counter, let me know.

The above mentioned amplifier has been used to amplify the input to the frequency counter. The frequency counter is now getting an input that it can read.

I have done an edit to the ALS medm screen and the PSL and AUX Y laser beat frequency is now readable.

  10892   Tue Jan 13 04:57:26 2015 ericqUpdateCDSAll FE diagnostics back to green

I was looking into the status of IPC communications in our realtime network, as Chris suggested that there may be more phase missing that I thought. However, the recent continual red indicators on a few of the models made it hard to tell if the problems were real or not. Thus, I set out to fix what I could, and have achieved full green lights in the CDS screen. 

This required:

  • Fixing the BLRMS block, as was found to be a problem in ELOG 9911 (There were just some hanging lines not doing anything)
  • Cleaning up one-sided RFM and SHMEM communications in C1SCY, C1TST, C1RFM and C1OAF

The frontend models have been svn'd. The BLRMs block has not, since its in a common cds space, and am not sure what the status of its use at the sites is...

  10891   Tue Jan 13 03:58:27 2015 JenneUpdateLSCTransitioned to ASDC MICH (PRMI and PRFPMI)

[Jenne, Diego, EricQ]

Hopefully there will be more later, but Chiara just went down (network? other?  Q is in there right now looking at it), so this is a so-far-tonight elog.

We have successfully transitioned MICH over from REFL33Q to ASDC in both the PRMI and PRFPMI configurations.  Next up is to start reducing the CARM offset.


Resetting the REFL demod phases

I have been unable to lock the PRMI for more than teeny blips since Thursday.  So, tonight I finally got it to lock with MICH on AS55Q and PRCL on REFL33I, and used that to set the demod phases. 

PRMI, AS55Q and REFL33I MICH PRCL
Input matrix 1*AS55Q 1*R33I
Gain -7 -0.022
DoF trigger POP22I; 80up, 2down POP22I; 80up, 2down
FM trigger FM 2,6,8; 100up, 2down; 0.3sec FM 2,6,9; 100up, 2down; 0.01sec
Normalization n/a n/a
Output matrix 0.5*BS, -0.2625*PRM 1*PRM
UGF servo n/a

n/a

Setting the demod phases, I used an oscillation of 100 cts to PRM, at 400.123 Hz.

REFL 33 demod phase started at 148deg, now 133.2deg.

REFL165 phase started at -105.53, now -172.

No signal in REFL55????  Time series and spectra both look just like noise.  Need to check alignment of beam on PD, or if cables unplugged!!

REFL11 phase started at 16.75, now 18.9deg.

Was then able to lock on REFL 33 I&Q, like normal.

PRMI, REFL33I&Q MICH PRCL
Input matrix 1*R33Q 1*R33I
Gain 2.5 -0.02
DoF trigger POP22I; 80up, 2down POP22I; 80up, 2down
FMs FM 4,5 on FM 4,5 on
FM trigger FM 2,6,8; 100up, 2down; 0.3sec FM 2,6,9; 100up, 2down; 0.01sec
Normalization n/a n/a
Output matrix 0.5*BS, -0.2625*PRM 1*PRM
UGF servo n/a

n/a



Transitioning PRMI from REFL33Q to ASDC

With the PRMI locked on REFL33 I&Q, I found that a MICH offset of -5 counts gives a minimum in ASDC.  From my earlier elog this evening (http://nodus.ligo.caltech.edu:8080/40m/10887), I expect the minimum to be at +1.4nm.  This is only one point though, so I don't know the calibration of the MICH offset yet (we should get this calib during the day by looking at MICH-only).  Anyhow, this informed which side was positive and negative relative to my Optickle plots, so I know that I wanted positive offset in the MICH servo.

I was able to comfortably hold lock at +20 counts.  Looking at a calibration line at 143.125 Hz, I determined that I wanted the matrix element for ASDC to be -0.05.  After I made that transition using ezcastep, I put the POPDC normalization in.  At the time, POPDC was about 151counts, so I put 1/151 in the POPDC->Mich matrix element.

So, here were the final lock parameters.  Note that in PRMI-only, you can acquire lock like this, and with a variety of MICH offsets:

PRMI, ASDC and REFL33I MICH PRCL
Input matrix -0.05 * ASDC 1*R33I
Gain 2.5 -0.02
DoF trigger POP22I; 80up, 2down POP22I; 80up, 2down
FMs FM 4,5 on FM 4,5 on
FM trigger FM 2,6,8; 100up, 2down; 0.3sec FM 2,6,9; 100up, 2down; 0.01sec
Normalization  0.0066 n/a
Output matrix 0.5*BS, -0.2625*PRM 1*PRM
UGF servo n/a

n/a


Locking PRMI part of PRFPMI

Since the PRMI has been fussy, I'm including a brief note on the PRMI settings when the arms are held with ALS off by roughly 3nm.  To get to this point, we just ran the usual carm_cm_up script, and didn't let it run anymore when it asked for confirmation that PRMI was locked.

PRFPMI, Arms=ALS, PRMI=REFL33I&Q MICH PRCL CARM DARM
Input matrix 1*R33Q 1*R33I -1*alsX+1*alsY 1*alsX+1*alsY
Gain 2.5 -0.012 8 8
DoF trigger POP22I; 80up, 2down POP22I; 80up, 2down n/a n/a
FMs FM 4,5 on FM 4,5 on FM 1,2,3,5,6 FM 1,2,3,5,6
FM trigger FM 2,6,8; 50up, 2down; 0.3sec FM 1,2,6,9; 50up, 2down; 0.01sec n/a n/a
Normalization n/a n/a n/a n/a
Output matrix 0.5*BS, -0.2625*PRM 1*PRM 1*etmx+1*etmy -1*etmx+1*etmy
UGF servo n/a set to 150Hz n/a n/a

With MICH offset of -30 counts, AS port is pretty bright.  ASDC dark offset is set to -475.4 by the LSCoffsets script. with MICH offset = 0, ASDC_OUT is around 300counts.  But, with MICH offset = -30, ASDC_OUT is about 525 counts.  So, I put that 525 counts into the ASDC filterbank offset (so it is now the dark offset + this extra offset), so the ASDC offset is currently around -1,000.  This makes the ASDC signal roughly zero when I am ready to transition MICH over to it.  In principle I should probably set it so the average is the same as the MICH offset, but the noise is so high relative to that offset, that it doesn't matter.

After this, we engaged the CARM and DARM UGF servos.  MICH was gain peaking, so I think we might want to turn that one on too, rather than my by-hand turning down the gain.

The transition has been successful 4 or 5 times with the arms held off resonance at 3nm.  Once, we reduced the CARM offset as low as 1.7 (and had to lower the MICH gain to 1.5), but we were still hearing a woomp-woomp sound.  Not sure what that was from.  At this point, Chiara died, so we lost lock.  After that, we re-acquired lock a few more times, but MC keeps losing it.  We are still able to make the MICH to ASDC transition though, which is good.

The transition won't work if the PRCL UGF servo is not on.  The gain multiplier goes from about 1.1 up to 2.4, so the PRCL gain is certainly changing through the transition.

Diego has written up scripts for the individual UGF servos (look for an elog from him separately), so now the carm_cm_up script goes as far as locking the PRMI on REFL33 I&Q, and then it starts to transition.  PRCL UGF is engaged, MICH offset is set to -30 counts, MICH is transitioned to ASDC, POP normalization engaged, CARM UGF servo turned on, and DARM UGF servo turned on.  There are "read"s in the script before each step, so you can stop whenever you like.

Here's the final configuration for the PRFPMI while the arms are held at 3nm, with MICH on ASDC (so after the transition):

PRFPMI, Arms=ALS, PRCL=REFL33I, MICH=ASDC MICH PRCL CARM DARM
Input matrix 0.27*ASDC 1*R33I -1*alsX+1*alsY 1*alsX+1*alsY
Gain 2.5 -0.015 8 8
DoF trigger POP22I; 80up, 2down POP22I; 80up, 2down n/a n/a
FMs FM 4,5 on FM 4,5 on FM 1,2,3,5,6 FM 1,2,3,5,6
FM trigger FM 2,6,8; 50up, 2down; 0.3sec FM 1,2,6,8,9; 50up, 2down; 0.01sec n/a n/a
Normalization 0.0042*POPDC n/a n/a n/a
Output matrix 0.5*BS, -0.2625*PRM 1*PRM 1*etmx+1*etmy -1*etmx+1*etmy
UGF servo n/a set to 150.001Hz set to 115.1Hz set to 110.1

 

The transition for MICH to ASDC has been successful with the arms held off resonance several times tonight. It's all part of the carm_cm_up script now. I think that if we hadn't lost about an hour of time and our momentum, we would have gotten farther.  I have high hopes for tomorrow night!

  10890   Tue Jan 13 03:47:46 2015 ericqUpdateCDSInstalled new ethernet driver on Chiara

Chiara threw another network hissy fit. Dmesg was spammed with a bunch of messages like eth0: link up appearing rapidly. 

Some googling indicated that this error message in conjuction with the very ethernet board and driver that Chiara had in use could be solved by updating with an appropriate driver from the manufacturer.

In essence, I followed steps 1-7 from here: http://ubuntuforums.org/showthread.php?t=1661489

So far, so good. We'll keep an eye out to see how it works...

  10889   Tue Jan 13 01:58:16 2015 ericqUpdateComputer Scripts / ProgramsCdsutils upgraded to 382

I've upgraded our cdsutils installation to v382; there have been some changes to pydv which will allow me to implement the auto y-scaling on our lockloss plots. 

After some brief testing, things seem to still work...

  10888   Tue Jan 13 01:11:51 2015 diegoUpdateLSCResponse of error signals to MICH EXC

For several MICH offsets, I measured the response of REFL33Q, ASDC and the ratio ASDC/POPDC to a MICH EXC. It appears that there is no frequency-dependent effect. The plots for MICH_OFFSET = 0.0 and 2.0 are slightly lower in magnitude: the reason is they were the first measurements done, and after that a little realignment of BS was necessary, so probably that is the reason.

 

 

Attachment 1: MICH_to_REFL33Q_ASDC_12Jan2015_1.pdf
MICH_to_REFL33Q_ASDC_12Jan2015_1.pdf
Attachment 2: MICH_to_REFL33Q_ASDC_12Jan2015_2.pdf
MICH_to_REFL33Q_ASDC_12Jan2015_2.pdf
Attachment 3: MICH_to_REFL33Q_ASDC_12Jan2015_3.pdf
MICH_to_REFL33Q_ASDC_12Jan2015_3.pdf
  10887   Tue Jan 13 00:42:15 2015 JenneUpdateLSCError signals for MICH with variable finesse technique

In order to know where we should try to make the transition from REFL##Q to ASDC for MICH, I did a quick Optickle simulation to see what the error signals will look like.

The idea is to try to lock the PRMI on a single REFL diode (ex. REFL33 I&Q) with some MICH offset, and then transition over to ASDC.  As soon as we have completed the transition, we can engage the normalization matrix to normalize ASDC by POPDC, and also increase the MICH offset if we want.  Unfortunately, we do not as yet have the ability in our model to independently normalize different error signals, and then blend them, so we have to turn on the normalization after we've transitioned.

Here is the situation for PRMI-only:

You can see that REFL33Q has a slightly wider range than REFL165Q.  It seems like we can perhaps try to make the transition around -15nm or so.  Note that the error signals are not quite symmetric about 0nm, so we can use that to help determine what + and - mean.  We expect that we need to add about 1nm offset to REFL33Q to get a true minimum in ASDC, so the sign of the digital offset that we need will tell us if there is a sign flip or not between the digital offset and this x-axis. 

After we get this to work (hopefully in the next hour or so....), we will want to try the same thing with the arms held off resonance. 

Usually we lock the PRMI at an offset of about 3nm:

However we could do it lower, perhaps around 1nm (which is where we currently are doing our CARM/DARM ALS->IR signals transitions):

At some point, we will arrive at 0nm CARM offset, when we'll want to transition back to RF signals (although probably we could jump straight to a 1f signal, not plotted):

The moral of the story here is that I'm not sure how we were ever successfully locking MICH on REFL165Q, unless my phase-setting in Optickle is way off.  Certainly it looks like we should be sticking with REFL33 for PRFPMI.  Also, since we have an offset in REFL33Q anyway (which we have seen and have commented on before), at 3nm CARM offset it looks like we could try to just do the jump without any extra digital offset.  Here's a zoom of the 3nm situation:

  10886   Mon Jan 12 18:11:25 2015 ericqUpdateASCTest Mass -> Transmon QPD TFs measured

We want to have some angular control of the arms during lock acquistion. 

In single arm lock, Diego and I shook the TMs and measured how the QPDs responded. (I would've liked to do a swept sine in DTT, but the user envelope function still isnt' working!)

For now, we can close simple loops with QPD sensor and ITM actuator, but, as Rana pointed out to Diego and me today, this will drive some amount of the angular cavity degree of freedom that the QPD doesn't sense. So, ideally, we want to come up with the right combination of ITM and ETM motion that lies entirely within the DoF that the QPD senses.

I created a rudimentary loop for Yarm yaw, was able to get ~20Hz for the upper UGF, a few mHz for the lower, but it was starting to leak into the length error signal. Further tweaking will be neccesary...

Attachment 1: Jan12_singleArmSensing.pdf
Jan12_singleArmSensing.pdf
Attachment 2: Jan12_singleArmSensing.xml.zip
  10885   Fri Jan 9 19:18:51 2015 JenneUpdatePSLPMC realigned

A few hours ago I tweaked up the alignment to the PMC.  It was really bad in pitch, and the transmission was down to about 0.711.

ELOG V3.1.3-